# Unique3D: High-Quality and Efficient 3D Mesh Generation from a Single Image

### v2

@@ Abstract
Figure 1: Gallery of Unique3D. High-fidelity and diverse textured mesh generated by Unique3D from single-view wild images within 30 seconds.
@@ 3 Method
In this section, we introduce our framework, i.e., Unique3D, for high-fidelity, efficient, and generalizable 3D mesh generation from a single wild image. Given an input image, we first generate four orthographic multi-view images with their corresponding normal maps from a multi-view diffusion model and a normal diffusion model. Then, we lift them to high-resolution space progressively (Sec 3.1). Given high-resolution multi-view RGB images and normal maps, we finally reconstruct high-quality 3D meshes with our instant and consistent mesh reconstruction algorithm ISOMER (Sec 3.2). An overview of our framework is depicted in Figure 2.
Figure 2: Pipeline of our Unique3D. Given a single wild image as input, we first generate four orthographic multi-view images from a multi-view diffusion model. Then, we progressively improve the resolution of generated multi-views through a multi-level upscale process. Given generated color images, we train a normal diffusion model to generate normal maps corresponding to multi-view images and utilize a similar strategy to lift it to high-resolution space. Finally, we reconstruct high-quality 3D meshes from high-resolution color images and normal maps with our instant and consistent mesh reconstruction algorithm ISOMER.
@@ 3.1 High-resolution Multi-view Generation
High-resolution Multi-view Image Generation. Instead of training from scratch, we start with the initialization of the pre-trained 2D diffusion model using the checkpoint of Stable Diffusion [53] and encode multi-view dependencies to fine-tune it to obtain a multi-view diffusion model that is able to generate four orthographic view images (256 resolution) from a single wild image. It is worth noting that the images generated in this step have relatively low resolution and suffer from multi-view inconsistency in out-of-the-domain data. This significantly limits the quality of recent works [23, 4, 5, 3, 51]. In contrast, we address the multi-view consistency issue during the reconstruction phase (Sec 3.2). Given the generated four orthographic view images, we then finetune a multi-view aware ControlNet [54] to improve the resolution of images. This model leverages the four collocated RGB images as control information to generate corresponding clearer and more precise multi-view results. It enhances the details and ameliorates unclear regions, leading the resolution of images from 256 to 512. Finally, we employ a single-view super-resolution model [55] to further upscale the image by a factor of four, achieving a resolution of 2048 that offers sharper edges and details without disrupting the multi-view consistency.
@@ 3.2 ISOMER: An Efficient Method for Direct Mesh Reconstruction
[…]
Although the diffusion process generates pseudo normal maps, these maps do not yield a real normal field which is irrotational. To address this, we introduce a random rotation to the normal map before integration. The process is repeated several times, and the mean value of these integrations is then utilized to calculate the depth, providing a reliable estimation. Subsequently, we map each pixel to its respective spatial location using the estimated depth, creating mesh models from both the front and back views of the object. The two models are seamlessly joined through Poisson reconstruction, which guarantees a smooth connection between them. Finally, we simplify them into 2000 fewer faces for our mesh initialization.
@@
ExplicitTarget Optimization for Multi-view Inconsistency and Geometric Refinement. Due to inherent inconsistencies in generated multi-view images from out-of-distribution (OOD) wild input, no solution can perfectly align with every viewpoint. After the above steps, we can only reconstruct a model that roughly matches the shape but lacks detail, falling short of our pursuit of high-quality mesh. Therefore, we cannot use the common method that minimizes differences in all views, which would lead to significant wave-pattern flaws, as shown in Figure 3-(a). To overcome this challenge, finding a more suitable optimization target becomes crucial. Under single-view supervision, although a complete model cannot be reconstructed, the mesh shape within the visible area of that view can meet the supervision requirements with highly detailed structures. Based on this, we propose a novel method that assigns a unique optimization target for each vertex to guide the optimization direction. In contrast to the conventional implicit use of multi-view images as optimization targets, we explicitly define the optimization target with better robustness. We call this explicit optimization target as ExplicitTarget and devise it as follows:
(ExplicitTarget). Let $P(v,i):(\mathbb{R}^{3},\mathbb{N}^{+})\rightarrow\mathbb{R}^{2}$ be the image space coordinates of a spatial point $v$ in view $i$, $Col(p,\mathcal{I}_{m}):(\mathbb{R}^{2},\mathbb{R}^{H\times W\times 3})\rightarrow\mathbb{R}^{3}$ denote the color of point $p$ in image $\mathcal{I}_{m}$, and $V_{M}(v,i):(\mathbb{N}^{+},\mathbb{N}^{+})\rightarrow\{0,1\}$ represent the visibility of vertex $v$ in mesh $M$ under view $i$. We compute the ExplicitTarget $ET$ of mesh $M$ as a mapping function from the set of vertices in $M$ to a set of colors:
$ET_{M}(\mathcal{I},\mathcal{I}_{m},v)=\begin{cases}\frac{\sum_{i\in\mathcal{I}}V_{M}(v,i)W_{M}(v,i)^{2}Col(P(v,i),\mathcal{I}_{m}^{(i)})}{\sum_{i\in\mathcal{I}}V_{M}(v,i)}&\text{, if }\sum_{i\in\mathcal{I}}V_{M}(v,\mathcal{I})>0\\ \mathbf{0}&\text{, otherwise, }\end{cases}$ (5)
where $\mathcal{I}$ refer to the set of sampling views, $\mathcal{I}_{m}$ are images corresponding to the views, and $W_{M}(v,i)=-\cos(N_{v}^{(M)},N_{i}^{(view)})$ is a weighting factor, where $N_{v}^{(M)}$ is the vertex normal of $v$ in mesh $M$, and $N_{i}^{(view)}$ is the view direction of view $i$.
In the function $ET_{M}(\mathcal{I},\mathcal{I}_{m})$, the result for vertex $v$ is computed as the weighted sum of supervised views, with weights determined by the square of cosine angles. This is because the projected area is directly proportional to the cosine value, and the prediction accuracy is also positively correlated with the cosine value. The object loss function for ExplicitTarget is defined as
@@ 4.1 Experimental Setting
Training Details: The entire training takes around 4 days on 8 NVIDIA RTX4090 GPUs. The primary level of multiview image generation uses $30k$ training iterations with a batch size of $1,024$. The second level of training involves $10k$ iterations with a batch size of $128$. Normal map prediction is trained for $10k$ iterations at a batch size of $128$. Additional training specifics are accessible in the Appendix.
@@ 4.2 Comparisons
Qualitative Comparison: To highlight the advantages of our methodology, we perform a comprehensive comparison with existing works, including CRM [4], one-2-3-45 [50], SyncDreamer [21], Wonder3D [3], OpenLRM [67], InstantMesh [5], and GRM [23]. For a fair quality comparison, we choose to present samples previously selected in the referenced papers, originating from Wonder3D [3], SyncDreamer [21], CRM [4], and InstantMesh [5]. The results are shown in Figure 3. Our results clearly surpass the existing works in both geometric and material quality, thereby emphasizing the benefits of our approach in achieving high resolution and intricate details in both geometry and material. In addition to the above overall quality comparison, we further show the comparison of the details in Figure 4, highlighting the advantage of our method in high resolution. The reconstruction process of ISOMER is completed in under $10$ seconds, while the entire procedure from the input image to high-precision mesh is accomplished in less than $30$ seconds on an RTX4090.
@@
One-2-3-45 [50] 13.19 0.7231 0.3795 0.7792 0.0246 0.4647 0.6025
OpenLRM [67] 14.73 0.7722 0.3294 0.8607 0.0229 0.4452 0.5907
SyncDreamer [21] 13.13 0.7262 0.4036 0.7828 0.0207 0.4220 0.5689
Wonder3D [3] 15.66 0.7898 0.3006 0.8612 0.0184 0.5199 0.6513
InstantMesh [5] 15.30 0.7809 0.3028 0.8998 0.0164 0.5141 0.6421
GRM [23] 14.50 0.7676 0.3332 0.8820 0.0161 0.4768 0.6178
CRM [4] 16.45 0.7965 0.2750 0.8936 0.0156 0.5379 0.6697
Ours 16.48 0.8007 0.2624 0.9096 0.0145 0.5538 0.6845
Quantitative Comparison: In line with previous work, we evaluate our results using the Google Scanned Objects (GSO) [69] dataset. We randomly select 30 objects and render frontal views at a resolution of $1024\times 1024$ as input for all methods. All generated mesh results are normalized to the bounding box $[-0.5,0.5]$ to ensure alignment. The geometric quality is assessed by calculating the distance to the ground truth mesh using metrics such as Chamfer Distance (CD), Volume IoU, and F-Score. Concurrently, we render $24$ views around the object, selecting one of $[0,15,30]$ for elevation angles and $8$ evenly distributed azimuth angles spanning a full 360-degree rotation. We employ PSNR, SSIM, LPIPS, and Clip-Similarity [68] to evaluate the visual quality. The results are presented in Table 1. As evidenced in table, both our geometric and material quality significantly outperform those of existing methods.
@@ 4.3 Ablation Study and Disscussion
Ablation Study: We analyze the importance of ExplicitTarget and expansion regularization in ISOMER. We compare samples with and without ExplicitTarget and Expansion Regularization in figure 5. We clearly show the improvement of ExplicitTarget for geometry and the necessity of expansion regularization for reconstruction. ExplicitTarget notably improves reconstruction results in challenging cases, while expansion regularization avoids some possible collapses.
@@ Appendix D ExplicitTarget algorithm
Appendix E Ablation Study on Mesh Initialization
@@
Appendix F User Study
@@
Appendix G Social Impact
@@
Appendix H Licenses for Used Assets

### v3

@@ Abstract
Figure 1: Gallery of Unique3D. High-fidelity and diverse textured mesh generated by Unique3D from single-view wild images within 30 seconds. Project Link.
@@ 3 Method
In this section, we introduce our framework, i.e., Unique3D, for high-fidelity, efficient, and generalizable 3D mesh generation from a single in-the-wild image. Given an input image, we first generate four orthographic multi-view images with their corresponding normal maps from a multi-view diffusion model and a normal diffusion model. Then, we lift them to high-resolution space progressively, (Sec 3.1). Given high-resolution multi-view RGB images and normal maps, we finally reconstruct high-quality 3D meshes with our instant and consistent mesh reconstruction algorithm ISOMER, (Sec 3.2). ISOMER directly handles the case where the global normal of the same vertex is inconsistent across viewpoints to enhance the consistency. An overview of our framework is depicted in Figure 2.
Figure 2: Pipeline of our Unique3D. Given a single in-the-wild image as input, we first generate four orthographic multi-view images from a multi-view diffusion model. Then, we progressively improve the resolution of generated multi-views through a multi-level upscale process. Given generated color images, we train a normal diffusion model to generate normal maps corresponding to multi-view images and utilize a similar strategy to lift it to high-resolution space. Finally, we reconstruct high-quality 3D meshes from high-resolution color images and normal maps with our instant and consistent mesh reconstruction algorithm ISOMER.
@@ 3.1 High-resolution Multi-view Generation
High-resolution Multi-view Image Generation. Instead of training from scratch, we start with the initialization of the pre-trained 2D diffusion model using the checkpoint of Stable Diffusion [53] and encode multi-view dependencies to fine-tune it to obtain a multi-view diffusion model that is able to generate four orthographic view images (256 resolution) from a single in-the-wild image. It is worth noting that the images generated in this step have relatively low resolution and suffer from multi-view inconsistency in out-of-the-domain data. This significantly limits the quality of recent works [23, 4, 5, 3, 51]. In contrast, we address the multi-view consistency issue during the reconstruction phase (Sec 3.2). Given the generated four orthographic view images, we then finetune a multi-view aware ControlNet [54] to improve the resolution of images. This model leverages the four collocated RGB images as control information to generate corresponding clearer and more precise multi-view results. It enhances the details and ameliorates unclear regions, leading the resolution of images from 256 to 512. Finally, we employ a single-view super-resolution model [55] to further upscale the image by a factor of four, achieving a resolution of 2048 that offers sharper edges and details without disrupting the multi-view consistency.
@@ 3.2 ISOMER: An Efficient Method for Direct Mesh Reconstruction
$d(i,j)=\sum_{t}=0^{i}n_{x}(t,j)$ (1)
where $n_{x}(t,j)$ is the normal vector of the $t$-th pixel in the $j$-th row. Although the diffusion process generates pseudo normal maps, these maps do not yield a real normal field which is irrotational. To address this, we introduce a random rotation to the normal map before integration. The process is repeated several times, and the mean value of these integrations is then utilized to calculate the depth, providing a reliable estimation. Subsequently, we map each pixel to its respective spatial location using the estimated depth, creating mesh models from both the front and back views of the object. The two models are seamlessly joined through Poisson reconstruction, which guarantees a smooth connection between them. Finally, we simplify them into $2000$ fewer faces for our mesh initialization.
@@
ExplicitTarget Optimization for Multi-view Inconsistency and Geometric Refinement. Due to inherent inconsistencies in generated multi-view images from out-of-distribution (OOD) in-the-wild input, no solution can perfectly align with every viewpoint. After the above steps, we can only reconstruct a model that roughly matches the shape but lacks detail, falling short of our pursuit of high-quality mesh. Therefore, we cannot use the common method that minimizes differences in all views, which would lead to significant wave-pattern flaws, as shown in Figure 5-(a). To overcome this challenge, finding a more suitable optimization target becomes crucial. Under single-view supervision, although a complete model cannot be reconstructed, the mesh shape within the visible area of that view can meet the supervision requirements with highly detailed structures. Based on this, we propose a novel method that assigns a unique optimization target for each vertex to guide the optimization direction. In contrast to the conventional implicit use of multi-view images as optimization targets, we explicitly define the optimization target with better robustness. We call this explicit optimization target as ExplicitTarget and devise it as follows:
(ExplicitTarget). Let $Avg(V,W)=\frac{\sum_{i}{V_{i}W_{i}}}{W_{i}}$ represent the weighted average function, and $V_{M}(v,i):(\mathbb{N}^{+},\mathbb{N}^{+})\rightarrow\{0,1\}$ represent the visibility of vertex $v$ in mesh $M$ under view $i$. $Col_{M}(v,i)$ Indicate the color of vertex v in viewpoint i. We compute the ExplicitTarget $ET$ of each vertex in mesh $M$ as
$ET_{M}(v)=\begin{cases}Avg(Col_{M}(v,i),V_{M}(v,i)W_{M}(v,i)^{2})&\text{, if }\sum_{i}V_{M}(v,\mathcal{i})>0\\ \mathbf{0}&\text{, otherwise, }\end{cases}$ (5)
[…]
In the function $ET_{M}(\mathcal{I},\mathcal{I}_{m})$, the predicted color of vertex $v$ is computed as the weighted sum of supervised views, with weights determined by the square of cosine angles. This is because the projected area is directly proportional to the cosine value, and the prediction accuracy is also positively correlated with the cosine value. The object loss function for ExplicitTarget is defined as
@@ 4.1 Experimental Setting
Training Details: The entire training takes around 4 days on 8 NVIDIA RTX4090 GPUs. The primary level of multiview image generation uses $30k$ training iterations with a batch size of $1,024$. The training of multi-view image upscaling involves $10k$ iterations with a batch size of $128$. Normal map prediction is trained for $10k$ iterations at a batch size of $128$. Additional training specifics are accessible in the Appendix.
@@ 4.2 Comparisons
Qualitative Comparison: To highlight the advantages of our methodology, we perform a comprehensive comparison with existing works, including CRM [4], one-2-3-45 [50], Wonder3D [3], OpenLRM [67], and InstantMesh [5]. For a fair quality comparison, we choose to present samples previously selected in the referenced papers, originating from Wonder3D [3], SyncDreamer [21], CRM [4], and InstantMesh [5]. The results are shown in Figure 3. Our results clearly surpass the existing works in both geometric and material quality, thereby emphasizing the benefits of our approach in achieving high resolution and intricate details in both geometry and material. In addition to the above overall quality comparison, we further show the comparison of the details in Figure 4, highlighting the advantage of our method in high resolution. The reconstruction process of ISOMER is completed in under $10$ seconds, while the entire procedure from the input image to high-precision mesh is accomplished in less than $30$ seconds on an RTX4090.
@@
One-2-3-45 16.1058 0.8874 0.1812 0.7782 0.0313 0.4142 0.5518
OpenLRM 18.0433 0.8957 0.1560 0.8416 0.0336 0.3947 0.5354
Wonder3D 18.0932 0.8995 0.1536 0.8535 0.0261 0.4663 0.6016
[…]
Unique3D w/o ET 20.0383 0.9199 0.1129 0.8675 0.0158 0.5320 0.6594
Wonder3D+ISOMER 18.6131 0.9026 0.1470 0.8621 0.0244 0.4743 0.6088
Quantitative Comparison: In line with previous work, we evaluate our results using the Google Scanned Objects (GSO) [69] dataset. We render frontal views at a resolution of $1024\times 1024$ with Blender EEVEE as input for all methods. All generated mesh results are normalized to the bounding box $[-0.5,0.5]$ to ensure alignment. The geometric quality is assessed by calculating the distance to the ground truth mesh using metrics such as Chamfer Distance (CD), Volume IoU, and F-Score. Concurrently, we render $24$ views around the object, selecting one of $[0,15,30]$ for elevation angles and $8$ evenly distributed azimuth angles spanning a full 360-degree rotation. We employ PSNR, SSIM, LPIPS, and Clip-Similarity [68] to evaluate the visual quality. The results are presented in Table 1. As evidenced in table, both our geometric and material quality outperform those of existing methods. We find that ISOMER can even be used to improve the consistency of other methods. For example, in Table 1, we replace Wonder3D’s reconstruction method with ISOMER, which is not only faster but also of higher quality.
@@ 4.3 Ablation Study and Disscussion
We analyze the importance of ExplicitTarget and expansion regularization in ISOMER. We compare samples with and without ExplicitTarget and Expansion Regularization in figure 5. We clearly show the improvement of ExplicitTarget for geometry and the necessity of expansion regularization for reconstruction. ExplicitTarget notably improves reconstruction results in challenging cases, while expansion regularization avoids some possible collapses.
Figure 6: Ablation on Colorize. We show a comparison of whether or not to apply ExplicitTarget in coloring, and we can see that the group that does not use ExplicitTarget has significant artifacts, as there is no precise consistency across multiple views.
Mirroring the approach used for geometry, we will include additional experimental results in Figure 6 that illustrate the impact of the Explicit Target method on texture quality. Without the Explicit Target, the results are obviously flawed.
Table 2: Quantitative comparison results for ablation on 100 random samples with random rotation on GSO dataset.
Method PSNR↑ SSIM↑ LPIPS↓ Clip-Sim↑ CD↓ Vol. IoU↑ F-Score↑
Unique3D 19.6744 0.9217 0.1101 0.8864 0.0118 0.5463 0.6833
Additionally, we added a new test with randomly rotated objects sampled from $azimuth\in U[-180,180],elevation\in U[-30,30]$ in Table 2 to test robustness in non-front-facing views. The test results show that Unique3D still performs well in this case, and even the geometry prediction is more accurate.
Figure 7: Ablation on Resolution. The visualization of the generated multi-views images at different stages is shown. Multi-level super-resolution does not change the general structure, but only improves the detail resolution, allowing the model to remain well-detailed.
We expand our study to include a qualitative comparison across various resolutions in order to demonstrate the differences between different resolutions in Figure 7. The results demonstrate the necessity of high resolution maps in generating high resolution meshes.
Figure 8: Challenging examples.
Challenging Examples: The majority of existing common samples are overly simplistic to effectively demonstrate the advantages of our study. Consequently, we select two complex samples: an object featuring detailed text and a photograph of a human, as shown in Figure 8. It is observable that our method exhibits exceptional mesh materials and geometry, even capable of sculpting geometric structures with textual detail. In the context of photographs, our reconstruction results are nearly on par with specialized image-to-character mesh generation methods.
@@ Appendix D ExplicitTarget algorithm
Appendix E More on Mesh Optimizations
Edge Collapse: This operation is used to avoid and heal defects in the mesh. It involves selecting an edge within a triangle and collapsing it to the other edge, effectively merging the two triangles into a single triangle. This process can help to eliminate narrow triangles that might be causing issues in the mesh, such as those that are too thin to accurately represent the surface they are approximating. Edge collapse can prevent the creation of topological artifacts and maintain the quality of the mesh. Edge Split: This is the opposite of edge collapse. In edge split, an edge that is longer than a specified maximum length is divided into two, creating new vertices at the midpoint of the edge. This operation is used to refine the mesh, ensuring that the local edge length is kept close to the optimal length. It helps to maintain the quality of the mesh by avoiding edges that are too long, which could lead to an inaccurate representation of the surface. Edge Flip: Edge flip is an operation that adjusts the connectivity of the mesh to improve its quality. It involves flipping an edge within a triangle to connect two non-adjacent vertices, effectively changing the triangulation of the mesh. This can help to maintain the degree of the vertices close to their optimal value, which is typically six for internal vertices (or four for boundary vertices). The goal of these operations is to improve the mesh quality while avoiding defects and ensuring that the mesh accurately represents the target geometry.
Appendix F Ablation Study on Mesh Initialization
@@
Appendix G User Study
@@
Appendix H Social Impact
@@
Appendix I Licenses for Used Assets